2m urea solution is diluted from 2Kg to 5Kg by addition of water. calculate molality of diluted solution.

Grade:

1 Answers

Sunil Kumar FP
9 years ago
molality is the number of moles of solute dissolved in 1 kg of solvent
Inchemistry, themolality,b(orm), of asolutionis defined as theamount of substance(inmol) of solute, divided by themass(inkg) of thesolvent, (notthe mass of the solution):

[b = \frac{n_{solute}}{m_{solvent}}]
now,
2=mole of urea/2
mole of urea=4

now new molality=4/5=.2 molal